I have a word problem as states Sara has a favorite print whose length is 3 inches longer than twice the width. If the area is 90 square inches find its dimensions?
Answer by checkley77(12844)

L=2W+3
LW=90
(2W+3)W=90
2W^2+3W-90=0
(2W+15)(W-6)=0
W-6=0
W=6 ANS. FOR THE WIDTH.
L=2*6+3
L=12+3=15 ANS. FOR THE LENGTH.
PROOF:
6*15=90
90=90